SEMI-EXPERIMENTAL EQUILIBRIUM STRUCTURES; RESULTS FROM DIFFERENT CALCULATIONS OF ALPHAS AND COMPARISON WITH THEORY

\underline{NORMAN C. CRAIG}, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; DAVID FELLER, Department of Chemistry, Washington State University, Pullman, WA 99164-4630; PETER GRONER, Department of Chemistry, University of Missouri-Kansas City, Kansas City, KS 64110-2499; DONALD C. MCKEAN, School of Chemistry, University of Edinburgh, EH9 3JJ, U.K.

Semi-experimental equilibrium structures for ethylene, butadiene, and 1,1-difluorocyclopropane have been determined from ground state rotational constants and alphas (vibration-rotation constants) computed with scaled harmonic force constants. These results are compared with optimized structures from high-level ab initio calculations and with structures obtained with alphas computed by Gaussian 03 without force constant scaling. A similar, comparative analysis is reported for the equilibrium structures for cyclopropene and 3,3-difluorocyclopropene. Caution is advised in using alphas from Gaussian. Also reported is the length of a pure sp^{2}-sp^{2} single bond in twisted (90^{∘}) butadiene as computed with high-level ab initio calculations.